OpenStudy (anonymous):

In the rhombus, m1 = 18x, m2 = x + y, and m3 = 30z. Find the value of the variables x,y, and z.

OpenStudy (eyust707):



OpenStudy (eyust707):



OpenStudy (eyust707):

m1 = m2 = m3

OpenStudy (eyust707):

m1 = 18x m2 = x + y 18x = x+y solve for y Now you know what y is in terms of x x + y = 30z plug in for y now you have z in terms of x 30z = 18x plug in for z now your only in terms of x solve for x it all falls apart

OpenStudy (anonymous):

makes no sense to me

OpenStudy (eyust707):

still need help?

OpenStudy (anonymous):

HELPP ! @eyust707